OpenStudy (anonymous):

F(X)=X+3,G(Y)=-6SQUARE ROOT OF Y-1 G(F(X))= -6 SQUARE ROOT OF X+3-1 WHAT IS THE DOMAIN OF THE FUNCTION G(F(X))

OpenStudy (anonymous):

So the equations are: \[f(x) = 5 - x\] \[g(y) = \sqrt{y}\] Right?

OpenStudy (anonymous):

YEAH

OpenStudy (anonymous):

f(x) = y \[g(f(x)) = \sqrt{5 - x}\]

OpenStudy (anonymous):

YEAH

OpenStudy (anonymous):

WOULD IT BE ALL NUMBER LESS THAN OR EQUAL TO 5

OpenStudy (anonymous):

Do you see the graph? http://www.wolframalpha.com/input/?i= √%285+-x%29

OpenStudy (anonymous):

Domain: \[x ≤ 5\]

OpenStudy (anonymous):

SO IM RIGHT?????/

OpenStudy (anonymous):

You got it! :)

OpenStudy (anonymous):

I gave you a medal, didn't I? :P

OpenStudy (anonymous):

YEAH I HAVE OTHERS PROBLEMS I HAVE CAN YOU HELP ME PLEASE

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

Let me restart my computer first... Should take like 1-2 mins!